OpenStudy (anonymous):

If f(x) = −4x − 2, find f(−3) 10 −14 12 −10

OpenStudy (tkhunny):

Please substitute x = -3 and, after a little algebra, display the result.

OpenStudy (anonymous):

-14?

OpenStudy (anonymous):

@tkhunny

OpenStudy (whpalmer4):

\[f(x) = -4x-2\] \[f(-3) = -4(-3)-2 = \]

OpenStudy (anonymous):

I got the same answer but it didnt take me that long lol

OpenStudy (anonymous):

I got -14

OpenStudy (whpalmer4):

if you got that same answer, you got the wrong answer

OpenStudy (kaederfds):

f(-3) = -4(-3)-2 =12-2 = 10

OpenStudy (whpalmer4):

positive * positive = positive positive * negative = negative negative * negative = positive

OpenStudy (anonymous):

so the answer is A?

OpenStudy (anonymous):